OpenStudy (samanthagreer):

A dentist selected 16 patients at random from a population of more than 1000 patients and asked them how many times the patient flosses per week. The dentist made a line plot of the results. What is the estimated mode of the number of times all patients floss per week? A. 5 B. 5.5 C. 6 D. 7

OpenStudy (samanthagreer):

@Great_And_Powerful

OpenStudy (anonymous):

sorry love, idk

OpenStudy (samanthagreer):

ok thanks how about you @Godlovesme

OpenStudy (godlovesme):

One sec Sam

OpenStudy (samanthagreer):

kk

OpenStudy (godlovesme):

So we have 7 five times , 6 three times, 5 four times and 1 once and 2 one time we don't need the zeros

OpenStudy (samanthagreer):

thanks

OpenStudy (godlovesme):

7 *5 = 35 6*3= 18 5*4= 20 2*1= 2 1*1=1 If u add all it will be 76 then divide by 16 Hope that's right :)

OpenStudy (godlovesme):

Oh yw :)